Recent Log Entries

Beam

Log Number: 4162430Logbook(s): Entry Makers: brukerBody: 

New cathode makes good beam at 130 kV.

Multiple things have happened since last night:
1. New photocathode in new puck
2. Singlewound MFX2I01
3. Shukui made a controlled access to do something to the laser window.

Will try to isolate the root cause as soon as we can. May need to transistion to PSS kicker certification first.

Plunger method: current vs. power

Log Number: 4162428Logbook(s): Entry Makers: brukerBody: 

Investigating behavior of plunger power supply based on anode current vs. laser power.
Picoammeter range set to 20 nA.

/cs/prohome/apps/s/StepNGraph/3-4/fileio/data/...
2023-07-13_131908: laser power vs. attenuator (Figure 1 top).
2023-07-13_132327: gun picoammeter vs. attenuator (Figure 1 bottom). Suspicious saturation.

Figure 2: Current divided by power. Less efficiency the more power we apply, which makes sense if the voltage sags under load, but the current is so low I'm not sure if it's plausible. It could even be SCL.

Klixon Check of High-Density Accelerator Magnets

Log Number: 4162427Logbook(s): Body: 

Lee Broeker, Kevin Banks and I verified that the klixon chains connected to the High-Density magnets throughout the accelerator indicate a fault at their respective latching chassis in the Service Buildings. Previous Interlock Checks, completed earlier in the SAD, of the box power supplies, proved that the latching chassis' would trip their respective box supplies.

iocin1mag restoration failed

Log Number: 4162426Logbook(s): Tags: Body: restorelib 2.6 Target restore date was The output from the myRestore execution was as follows: <MyAPI>: Channel lookup fault [MBH0I06AH.HVAL]

COMMENT ON SRF Vacuum Control General Maintenance

Log Number: 4162422Logbook(s): Body: 

Valve cables for VBV2L03B and VBV2L04A have been re-terminated and a new jumper installed. Valve now cycle and are open.

Valve cables for VBV2L23B and VBV2L24A have been re-routed and are now able to cycle valves.

VBV2L22B and VBV2L23A valve cables will be re-routed after new air lines have been ran for them.

COMMENT ON 2L23 Repair

Log Number: 4162421Logbook(s): Body: 

Insulating vacuum leak check good. Minimal detectable leak 1e-8.
Helium circuit leak check. Detected a 2e-7 leak with circuit at 10psi using 3atm helium line. This cryomodule was delivered
to the linac with a -7 leak in the helium circuit. Twistorr insulating vacuum pump shouldn't have any problem maintaining at 4k. Return header feed through leak was sized at 3e-4.

Cryomodule ready for 4k cool down Monday.

MFX2I01 now configured as single-wound

Log Number: 4162415Logbook(s): Body: 

This morning, Sarin and I after performing LOTO of channel 7 in IN02B10 (MFX2I01), swapped the wires of the upstream coil at the terminal block. After the work on the wiring was completed, LOTO was removed and we were able to verify the single-wound configuration with a gauss meter.

We have photographs and will post them later.

Philip, Delk

COMMENT ON Modify LCW Piping

Log Number: 4162412Logbook(s): Body: 

Pump PMs have been completed for all LCW pumps, with the exception of the South Linac LCW pumps. PM will be completed for the SL before work begins on that system. I will update this Atlis when that PM is completed. Work will be begin on Tuesday July 18th starting with the West Arc LCW system.

Follow-up Re: NL dropped to LD

Log Number: 4162411Logbook(s): Body: 

Investigated and found the system A interface chassis for RF zones 17-21 had lost power do to bad connection at power connector. Repaired connection system powered up as designed, performed "wiggle" test all is good.

CTF Main C1 Restart

Log Number: 4162410Logbook(s): Entry Makers: rangharaBody: 

Turning ON main compressor C1 to get the plant ready for VTA ops.

CTF Expander 1 Startup and VTA Bayonet Can Flow

Log Number: 4162408Logbook(s): Entry Makers: rangharaBody: 

establishing flow back through the VTA bayonet can by EV27910. And also turning on the CB1 Expander. This is all part of getting CTF ops back to normalcy to support SRF operations.

NL dropped to LD

Log Number: 4162407Logbook(s): Entry Makers: mockBody: 

Earlier today the NL dropped to LD. Not sure of the exact cause, alarms showed RF 1L22 power supply unsafe but that is still locked out at breaker. A simple key flip brought it back to RA. Spoke with SSG.

The NL dropped to LD again. This time showing A chain RF waveguide pressure and off/safe are red. As well as A chain Off/safe on DOG magnets. Contacted SSG to investigate.

ODH Alarms

Log Number: 4162406Logbook(s): Entry Makers: dchriBody: 

Received 2 ODH alarms around 11:00am. Both Alarms were about 10min apart from one another. Looking at ODH summary screen and checking ODH levels. Nothing seems to be wrong and ODH levels are nominal.

Will continue to monitor the system and check for any possible ODH situations as they arise

HALL A SEE BPMs Troubleshoot - Chicane Line (IPM1P02A and IPM1P02B) BPMs - COMPLETE

Log Number: 4162400Logbook(s): Entry Makers: sethgBody: 

Starting yesterday I began troubleshooting the Chicane line BPMs in HALL A. I had found they were failing Alpha Calibration via Bad X & Y fit. When utilizing the Twiddle Tunnel for troubleshooting I found:

IPM1P02A - Y OSC On X- Channel BAD and X OSC on Y- Channel BAD

IPM1P0SB - Y OSC On X+ Channel BAD and X OSC on Y- Channel BAD

I also utilized a signal generator to inject signals into the respective IF Cards for these BPMs and found that the IF Cards were functioning properly.

iocw107rfsep1 restored

Log Number: 4162399Logbook(s): Tags: Body: restorelib 2.6 Target restore date was 2023-07-13 10:12:27 The output from the myRestore execution was as follows: iocw107rfsep1 Pause: 30 second(s) Execute: /cs/opshome/burt/REBOOT/Rf12Sep/myRestore-ical.csh 1 896 (success) Rf12Sep_W1P1234@2023-07-13 10:12 166(166) restored system@2023-07-13 10:12 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocw107rfsep1 (success) Procedure return status: Success

Pages